What's the solution to -3+8x=-x?
Found 2 solutions by gwendolyn, cleomenius:
Answer by gwendolyn(128)

First, we add x to both sides of the equation, so we can combine the two terms with 'x' in them.
-3+8x=-x
-3+8x+x=-x+x
-3+9x=0
Then, we can add 3 to both sides of the equation:
-3+9x+3=0+3
9x=3
Finally, we divide both sides of the equation by 9 to isolate the variable:
9x/9=3/9
x=1/3
So, x is equal to 1/3

Answer by cleomenius(959)

First, gather your like terms on the same side.
-3 + 8x = -x
7x = 3
x = 3/7
To check, 8(3/7 ) = 24/7
-21/7 + 24/7 = 3/7
This dose check.
Cleomenius.